What is the domain of the problems g(t) = sin(e to the negative t power) and the domain of g(t)= the square root of (1-2 to the t power)
Answer by stanbon(75887)

What is the domain of the problems
g(t) = sin(e to the negative t power)
---
g(t) = sin(e^-t)
-----
Comment: That doesn't look right. Are you sure that is the problem?
If that is right, the domain is "All Real Numbers".
==========================================================================
Domain of g(t)= the square root of (1-2 to the t power)
g(t) = sqrt[1 - 2^t]
---
1-2^t must be >= 0
2^t <= 1
t*log(2) <= log1
t <= 0/log(2)
t <= 0
-------------------------
So the domain is "All Real Numbers <= 0
============================================
Cheers,
Stan H.
============================================